OpenStudy (anonymous):

write an equation of the line that passes through the given point and is parallel to the graph of the given equation. (0, -5); y = 9x

OpenStudy (anonymous):

y = 9x + c makes it parallel then: -5 = 9(0) + c therefore c = -5 so y = 9x-5